# Union: A Trust-minimized Bridge for Bitcoin

### v1

@@ Abstract
We present Union, a trust-minimized bridge protocol that enables secure transfer of BTC between Bitcoin and a secondary blockchain. The growing ecosystem of blockchain systems built around Bitcoin has created a pressing need for secure and efficient bridges to transfer BTC between networks while preserving Bitcoin’s security guarantees. Union employs a multi-party variant of BitVMX, an optimistic proving system on Bitcoin, to create a bridge that operates securely under the assumption that at least one participant remains honest. This $1$-of-$n$ honest approach is strikingly different from the conventional honest-majority assumption adopted by practically all federated systems. The protocol introduces several innovations: a packet-based architecture that allows security bonds to be reused for multiple bridge operations, improving capital efficiency; a system of enablers to manage functionaries participation and to enforce penalties; a flexible light client framework adaptable to various blockchain architectures; and an efficient stop watch mechanism to optimize time-lock management. Union is a practical and scalable solution for Bitcoin interoperability that maintains strong security guarantees and minimizes trust assumptions.
@@ 1 Introduction
Since then, various approaches such as BitVM2 [9], BitVMX [7], SNARKnado [11], and BitSNARK [5] have emerged. As the universe of Bitcoin verification systems expanded, the next logical phase began. Building on these verifiers, several new bridging systems have been proposed. These advancements have led to the development of the BitVM bridge [9], Strata [12], Grail [5], and Clementine [4]. This paper contributes to this new phase of bridging systems with Union, a trust-minimized bridge protocol that enables the secure transfer of BTC between Bitcoin and secondary blockchain systems.
@@ 2.1 Protocol assumptions
  • Secondary blockchain ($S$): the protocol connects Bitcoin to an independent secondary blockchain system that we call $S$. This blockchain system $S$ maintains a globally replicated database [10].
@@
The Union protocol operates under the following assumptions:
@@ 3.2 Union bridge protocol overview
The Union protocol enables trust-minimized transfer of BTC between Bitcoin and a secondary blockchain system $S$. The protocol is operated by a set of functionaries who collectively manage a multi-party BitVMX setup on Bitcoin to secure the locked funds. Every peg-in locks some BTC inside a UTXO, each of which is secured by its own BitVMX setup. As we will see, separate BitVMX setups can share and reuse some components of the transaction graph. For now, we note that each peg-in creates a VMXO (a BitVMX secured UTXO). And ultimately, each peg-out will be sourced from a pre-existing VMXO.
The protocol operates under the assumption that at least one of $N$ functionaries is honest. When users want to transfer BTC to system $S$ (peg-in), they lock their funds in the BitVMX setup and receive a wrapped version of BTC on $S$. For transfers back to Bitcoin (peg-out), any functionary can act as a prover by advancing the BTC to the user and then recovering these funds from the BitVMX setup by proving the validity of the peg-out operation on $S$ and the transfer of the correct amount to the user on Bitcoin. The remaining functionaries act as verifiers in this process, ensuring the correctness of the proof. This arrangement allows any functionary to facilitate transfers while being verified by their peers, creating a flexible and secure bridge between the two blockchain systems.
@@ 3.3 Cross-chain validation
For the Union protocol to transfer BTC between Bitcoin and system $S$ in both directions, it is crucial to prove that specific events have occurred on one blockchain. This cross-chain validation is achieved through light-client implementations. In the case of peg-ins, we need to prove that a valid deposit transaction exists on Bitcoin. This is possible through $S$’s consensus because, as assumed in Section 2, $S$ implements a Bitcoin relay, and thus has the ability to validate a Bitcoin transaction inclusion. In the case of peg-outs, we need to prove on Bitcoin that a valid matching peg-out transaction exists on $S$. The Union protocol assumes that it is possible to implement a light client for $S$, which can then be validated using BitVMX.
@@ 3.4 Peg-in process
The peg-in process, which enables users to transfer BTC from Bitcoin to system $S$, uses the same multi-party BitVMX scheme depicted in Figure 2. The user deposits BTC using the Locking transaction shown in the diagram, and this triggers the unlocking of the wrapped BTC on $S$. Peg-ins are a permissioned process as the multi-party BitVMX setup requires all the functionaries to sign the transactions and delete their private keys. The system $S$ accepts a deposit transaction as a valid peg-in if it is signed by all the functioanries and confirmed on Bitcoin. This is in line with the 1-of-$n$ honest assumption of the Union protocol. The peg-in process consists of the following steps:
@@
    All functionaries collectively sign the deposit transaction. Since the input UTXOs belong to the user, this collective signing does not affect their spending. The collective signature is crucial as it serves to validate the deposit as a legitimate peg-in transaction on system $S$. This requirement ensures that the peg-in process also operates under the 1-of-$n$ honest assumption, as all functioanries must approve the deposit.
@@
    Once the deposit transaction is broadcast, and receives a required number of confirmations, any participant (functionary, user, or external observer) can inform system $S$ about it.
@@
    System $S$ then validates the transaction’s inclusion in Bitcoin and the functionaries’ signatures, after which it releases the corresponding amount of wrapped BTC to the user’s address on system $S$. This process is guaranteed by one of the protocol’s core assumptions: system $S$’s ability to validate Bitcoin transaction inclusion.
@@ 3.5 Peg-out process
The peg-out process, which enables users to transfer wrapped BTC from $S$ back to Bitcoin, operates as follows:
@@
    The user creates a peg-out transaction in system $S$, sending their wrapped BTC to a special peg-out address. Peg-out amounts are restricted to the same set of predetermined values as peg-ins.
@@
    The peg-out transaction is then uniquely linked to one of the UTXOs from a previous peg-in secured by a BitVMX setup. This linking can be handled either by the functionary set or automatically by system $S$.
@@
    After a required number of confirmations in system $S$, a functionary acting as the operator advances (or fronts) the BTC to the user on Bitcoin. As discussed later, the amount fronted to the user may be reduced as a fee. At this point, the user has effectively completed their peg-out.
@@
    After a number of confirmations on Bitcoin, the operator submits the Bitcoin transaction that advanced the funds to the user to system $S$ for validation. This generates a proof of peg-out on system $S$.
@@ 3.6 Security deposits
The protocol mandates that functionaries maintain security deposits to ensure proper behavior when acting as operators or verifiers. To keep the Union protocol agnostic about the particulars of the secondary chain $S$, security deposits are held on Bitcoin. Specific implementations of the protocol can choose to move some or all security bonds to $S$. These deposits serve multiple purposes. They incentivize honest behavior by making malicious actions economically costly. In addition, they allow honest functionaries to recover the costs incurred while proving their correct assertions when challenged to do so. For functionaries who act honestly, the only cost of operating the bridge should be the opportunity cost of their locked security deposits. We can distinguish between two types of costs in operating the bridge. First, there are the operational costs, which are the regular expenses incurred during normal bridge operations. These include transaction fees for standard peg-in and peg-out operations, infrastructure and maintenance costs, and administrative overhead. These costs should be covered by the fees paid by users for using the bridge services. Second, there are the security enforcement costs, which arise when dealing with malicious behavior. These include expenses related to executing the challenge-response protocol, conducting dispute resolution transactions, and performing emergency protocol operations. These costs should be covered by slashing the security deposits of the malicious actors involved. For the purpose of calculating security deposits, we assume that operational costs are covered by user fees, and focus on ensuring that security deposits are sufficient to cover all potential security enforcement costs.
@@ 5 Conclusions
In this paper, we present Union, a novel trust-minimized bridge protocol that enables secure transfer of BTC between Bitcoin and secondary blockchain systems. Union leverages BitVMX, an optimistic proving system on Bitcoin, in a multi-party setting to create a bridge that operates securely under the assumption that at least one participant remains honest. The protocol introduces several key innovations:

### v2

# Union: A Trust-minimized Bridge for Rootstock

@@ Abstract
We present Union, a trust-minimized bridge protocol that enables secure transfer of BTC between Bitcoin and Rootstock. The growing ecosystem of blockchain systems built around Bitcoin has created a pressing need for secure and efficient bridges to transfer BTC between networks while preserving Bitcoin’s security guarantees. Union employs a multi-party variant of BitVMX, an optimistic proving system on Bitcoin, to create a bridge that operates securely under the assumption that at least one participant remains honest. This $1$-of-$n$ honest approach is strikingly different from the conventional honest-majority assumption adopted by practically all federated systems. The protocol introduces several innovations: a packet-based architecture that allows security bonds to be reused for multiple bridge operations, improving capital efficiency; a system of enablers to manage functionaries participation and to enforce penalties; a flexible light client framework adaptable to various blockchain architectures; and an efficient stop watch mechanism to optimize time-lock management. Union is a practical and scalable solution for Bitcoin interoperability that maintains strong security guarantees and minimizes trust assumptions.
@@ 1 Introduction
Since then, various approaches such as BitVM2 [10], BitVMX [7], SNARKnado [11], and BitSNARK [5] have emerged. As the universe of Bitcoin verification systems expanded, the next logical phase began. Building on these verifiers, several new bridging systems have been proposed. These advancements have led to the development of the BitVM bridge [10], Strata [12], Grail [5], and Clementine [4]. This paper contributes to this new phase of bridging systems with Union, a trust-minimized bridge protocol that enables the secure transfer of BTC between Bitcoin and Rootstock [8], the longest running and leading Bitcoin sidechain.
@@ 2.1 Protocol assumptions
  • Rootstock: the protocol connects Bitcoin to Rootstock.
@@
The Union protocol is a bridge between Bitcoin and Rootstock. Nevertheless, the Union protocol can be generalized to any secondary blockchain system $S$ under the following assumptions:
@@ 3.2 Union bridge protocol overview
The Union protocol enables trust-minimized transfer of BTC between Bitcoin and Rootstock. The protocol is operated by a set of functionaries who collectively manage a multi-party BitVMX setup on Bitcoin to secure the locked funds. Every peg-in locks some BTC inside a UTXO, each of which is secured by its own BitVMX setup. As we will see, separate BitVMX setups can share and reuse some components of the transaction graph. For now, we note that each peg-in creates a VMXO (a BitVMX secured UTXO). And ultimately, each peg-out will be sourced from a pre-existing VMXO.
The protocol operates under the assumption that at least one of $N$ functionaries is honest. When users want to transfer BTC to Rootstock (peg-in), they lock their funds in the BitVMX setup and receive a wrapped version of BTC on Rootstock. For transfers back to Bitcoin (peg-out), any functionary can act as a prover by advancing the BTC to the user and then recovering these funds from the BitVMX setup by proving the validity of the peg-out operation on Rootstock and the transfer of the correct amount to the user on Bitcoin. The remaining functionaries act as verifiers in this process, ensuring the correctness of the proof. This arrangement allows any functionary to facilitate transfers while being verified by their peers, creating a flexible and secure bridge between the two blockchain systems.
@@ 3.3 Cross-chain validation
For the Union protocol to transfer BTC between Bitcoin and a generic secondary blockchain system $S$ (e.g., Rootstock), it is crucial to prove in one blockchain that specific events have occurred on the other blockchain. This cross-chain validation is achieved through light-client implementations. In the case of peg-ins, we need to prove that a valid deposit transaction exists on Bitcoin. This is possible through $S$’s consensus because, as assumed in Section 2, $S$ implements a Bitcoin relay, and thus has the ability to validate a Bitcoin transaction inclusion. In the case of peg-outs, we need to prove on Bitcoin that a valid matching peg-out transaction exists on $S$. The Union protocol assumes that it is possible to implement a light client for $S$, which can then be validated using BitVMX. Both of these assumptions are valid for the specific case of the Rootstock sidechain.
@@ 3.4 Peg-in process
The peg-in process, which enables users to transfer BTC from Bitcoin to Rootstock, uses the same multi-party BitVMX scheme depicted in Figure 2. The user deposits BTC using the Locking transaction shown in the diagram, and this triggers the unlocking of the wrapped BTC on Rootstock. Peg-ins are a permissioned process as the multi-party BitVMX setup requires all the functionaries to sign the transactions and delete their private keys. Rootstock accepts a deposit transaction as a valid peg-in if it is signed by all the functioanries and confirmed on Bitcoin. This is in line with the 1-of-$n$ honest assumption of the Union protocol. The peg-in process consists of the following steps:
@@
    All functionaries collectively sign the deposit transaction. Since the input UTXOs belong to the user, this collective signing does not affect their spending. The collective signature is crucial as it serves to validate the deposit as a legitimate peg-in transaction on Rootstock. This requirement ensures that the peg-in process also operates under the 1-of-$n$ honest assumption, as all functioanries must approve the deposit.
@@
    Once the deposit transaction is broadcast, and receives a required number of confirmations, any participant (functionary, user, or external observer) can inform Rootstock about it.
@@
    Rootstock then validates the transaction’s inclusion in Bitcoin and the functionaries’ signatures, after which it releases the corresponding amount of wrapped BTC to the user’s address on Rootstock. This process is guaranteed by one of the protocol’s core assumptions: Rootstock’s ability to validate Bitcoin transaction inclusion.
@@ 3.5 Peg-out process
The peg-out process, which enables users to transfer wrapped BTC from Rootstock back to Bitcoin, operates as follows:
@@
    The user creates a peg-out transaction in Rootstock, sending their wrapped BTC to a special peg-out address. Peg-out amounts are restricted to the same set of predetermined values as peg-ins.
@@
    The peg-out transaction is then uniquely linked to one of the UTXOs from a previous peg-in secured by a BitVMX setup. This linking can be handled either by the functionary set or automatically by Rootstock.
@@
    After a required number of confirmations in Rootstock, a functionary acting as the operator advances (or fronts) the BTC to the user on Bitcoin. As discussed later, the amount fronted to the user may be reduced as a fee. At this point, the user has effectively completed their peg-out.
@@
    After a number of confirmations on Bitcoin, the operator submits the Bitcoin transaction that advanced the funds to the user to Rootstock for validation. This generates a proof of peg-out on Rootstock.
@@ 3.6 Security deposits
The protocol mandates that functionaries maintain security deposits to ensure proper behavior when acting as operators or verifiers. These deposits are held in Bitcoin, although could also be held on Rootstock (or a generic secondary blockchain system). These deposits serve multiple purposes. They incentivize honest behavior by making malicious actions economically costly. In addition, they allow honest functionaries to recover the costs incurred while proving their correct assertions when challenged to do so. For functionaries who act honestly, the only cost of operating the bridge should be the opportunity cost of their locked security deposits. We can distinguish between two types of costs in operating the bridge. First, there are the operational costs, which are the regular expenses incurred during normal bridge operations. These include transaction fees for standard peg-in and peg-out operations, infrastructure and maintenance costs, and administrative overhead. These costs should be covered by the fees paid by users for using the bridge services. Second, there are the security enforcement costs, which arise when dealing with malicious behavior. These include expenses related to executing the challenge-response protocol, conducting dispute resolution transactions, and performing emergency protocol operations. These costs should be covered by slashing the security deposits of the malicious actors involved. For the purpose of calculating security deposits, we assume that operational costs are covered by user fees, and focus on ensuring that security deposits are sufficient to cover all potential security enforcement costs.
@@ 5 Conclusions
In this paper, we present Union, a novel trust-minimized bridge protocol that enables secure transfer of BTC between Bitcoin and Rootstock. Union leverages BitVMX, an optimistic proving system on Bitcoin, in a multi-party setting to create a bridge that operates securely under the assumption that at least one participant remains honest. The protocol introduces several key innovations: